OpenStudy (anonymous):

how do i implicitly differentiate (x^2)(y')

OpenStudy (amistre64):

product rule

OpenStudy (anonymous):

i know that, im confused on what the terms would be because of the y'

OpenStudy (amistre64):

y' is just another function in its own right: D[(x^2)(y')] = x'^2 y' + x^2 y'' = 2x y' + x^2 y''

OpenStudy (amistre64):

is this a part of a bigger problem?

OpenStudy (anonymous):

yeah it is, i think i see this part now though....could the second part be: x^2 yy' ? or is that not the same

OpenStudy (amistre64):

no, that is not the same; but i could have better answer prolly if I knew the problem from the start

OpenStudy (anonymous):

haha, well the full question is 9x3 + x2y − xy3 = 8

OpenStudy (amistre64):

\[9x^3 + x^2y − xy^3 = 8\] \[D[9x^3 + x^2y − xy^3 = 8]\] \[D[9x^3] + D[x^2y] − D[xy^3] = D[8]\] \[27x^2 + 2xy+x^2Y − (y^3+3xy^2Y) = 0\] \[27x^2 + 2xy+x^2Y − y^3-3xy^2Y = 0\] \[ x^2Y -3xy^2Y = -27x^2-2xy+ y^3\] \[ Y(x^2 -3xy^2) = -27x^2-2xy+ y^3\] \[ Y = \frac{-27x^2-2xy+ y^3}{x^2 -3xy^2}\] I used Y for y' just so it stands out, but is this good so far?

OpenStudy (anonymous):

yep, i actually can see that..i think, the problem i am working on breaks it into sections also so this is real helpful, ive been doing the problem and i have the same thing as you so far

OpenStudy (amistre64):

and you want to derive it once more right?

OpenStudy (anonymous):

nope, thats it! thanks!

OpenStudy (amistre64):

cool :)